# thanks: M.C. and D.P. contributed equally to this work

### v1

@@ II Regular black holes
Figure 1: Evolution of the temperatures (normalized to the temperature of Schwarzschild black holes, $T_{\text{Sch}}=1/8\pi M$) as a function of the regularizing parameter $\ell$ (normalized by the event horizon radius $r_{H}$) for the four regular black holes studied in the work: the Bardeen (blue solid curve, Sec. II.1), Hayward (red dashed curve, Sec. II.2), Ghosh-Culetu-Simpson-Visser (green dotted curve, Sec. II.3), and Hayward-like (magenta dash-dotted curve) regular space-times. In the first three cases the temperature is a monotonically decreasing function of $\ell$, whereas in the last case it increases monotonically with $\ell$ after an initial dip, with important phenomenological consequences resulting from these different behaviours. Note that the range of allowed values of $\ell/r_{H}$ is different for the four regular black holes.
@@
A final caveat is in order before discussing the RBH metrics we consider. The latter are all regular in the sense of having finite curvature invariants $R\equiv g^{\mu\nu}R_{\mu\nu}$, $R_{\mu\nu}R^{\mu\nu}$, and ${\cal K}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. However, a more stringent criterion for regularity is that of geodesic completeness, which does not necessarily imply finiteness of curvature invariants and viceversa. A number of “popular” RBHs have indeed been shown to have finite curvature invariants but to be geodesically incomplete Zhou and Modesto (2023). This includes the well-known Bardeen and Hayward RBHs, which are among the ones we shall consider here. However, given the significant interest in these metrics, the fact that they are widely taken as prototypes for RBHs, and our phenomenological goal of going beyond Schwarzschild PBHs, we will take these space-times into consideration, while cautioning the reader about the above issues, and therefore that these metrics should be considered nothing more than phenomenological toy models at this stage. Note, in addition, that the stability of RBHs featuring inner horizons is currently a matter of debate in the literature Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023).
@@ II.1 Bardeen black hole
The Bardeen BH is easily one of the best known RBHs, and one of the first ones to ever have been proposed Bardeen (1968). It is characterized by the following metric function:  [^2] We note that in all the metrics considered, the parameter $M$ appearing in the metric function can always be unambiguously identified with the BH mass (either the Komar, ADM, Misner-Sharp-Hernandez, or Brown-York mass). This is important as the later constraints on $f_{\text{pbh}}$ as a function of $M_{\text{pbh}}$ will identify the latter with $M$.
@@ II.2 Hayward black hole
Another widely known RBH space-time is the Hayward RBH Hayward (2006), characterized by the following metric function:
@@
If expressed in terms of BH mass $M$, the regularizing parameter for the Hayward BH is subject to the same limit as that of the Bardeen BH, i.e. $\ell\leq\sqrt{16/27}\,,M$. On the other hand, if expressed in terms of the more physically motivated horizon radius, the limit is instead $\ell\lesssim 0.57\,r_{H}$. We note that the Schwarzschild metric is function is recovered in the $\ell\to 0$ limit.
@@
Just as the Bardeen RBH possesses a dS core, so does the Hayward RBH. Indeed, introducing a dS core characterized by a (positive) cosmological constant $\Lambda=3/\ell^{2}$ in order to prevent the central singularity was precisely the original justification for the Hayward BH which, just like its Bardeen counterpart, was introduced on purely phenomenological grounds. Nevertheless, potential theoretical origins for the Hayward BH have been investigated, and range from corrections to the equation of state of matter at high density Sakharov (1966); Gliner (1966), finite density and finite curvature proposals Markov (1982, 1987); Mukhanov and Brandenberger (1992), theories of non-linear electrodynamics Kumar et al. (2020b); Kruglov (2021), and more generally as the result of corrections due to quantum gravity Addazi et al. (2022); Alves Batista et al. (2023). Just as with the Bardeen RBH, we shall here treat the Hayward RBH as a model-agnostic phenomenological toy model for a singularity-free space-time.
II.3 Ghosh-Culetu-Simpson-Visser black hole
@@
The regular space-times considered so far feature dS cores, which in itself is a very common feature of several RBH metrics. Nevertheless, another interesting phenomenological possibility consists in considering “hollow” RBHs in which the central singularity is replaced by an asymptotically Minkowski core, where the associated energy density and associated pressure asymptote to zero. This is quite unlike the case of the dS core where the energy density asymptotes to a finite value associated to a positive cosmological constant, and the pressure asymptotes to an equal but opposite value. Possible theoretical/mathematical motivations for considering RBHs with Minkowski cores include the fact that the vanishing energy density can significantly simplify the physics in the deep core, whereas the otherwise messy solutions to polynomical equations (which often cannot be written down in closed form) can be traded for arguably more elegant special functions, resulting in the space-time being more tractable. Our physical motivation in considering this class of BHs is instead to broaden the range of physical properties and phenomenological implications of PRBHs, going beyond the dS core RBHs studied thus far.
@@
With this in mind, we consider a RBH featuring a Minkowski core, independently studied by Ghosh Ghosh (2015), Culetu Culetu (2015), as well as Simpson and Visser Simpson and Visser (2019). Although such a RBH does not have any particular name associated to it in the literature, here we shall conform to the name introduced in Ref. Vagnozzi et al. (2023), referring to it as GCSV BH (from the initials of the four authors above). This space-time is characterized by the following metric function:
$f_{\text{GCSV}}(r)=1-\frac{2M}{r}\exp\left(-\frac{\ell}{r}\right)\,.$ (10)
@@
The horizon radius $r_{H}$, for which a closed form expression is not available in the Bardeen and Hayward cases, is given by:
@@
or, alternatively, $0\leq\ell<r_{H}$. While the GCSV BH was original introduced purely on phenomenological/mathematical grounds, it was shown in Refs. Kumar and Ghosh (2021); Singh et al. (2022) that such a space-time can emerge within the context of GR coupled to a specific non-linear electrodynamics source. In this case, denoting by $g$ the non-linear electrodynamics coupling constant/charge, the regularizing parameter $\ell$ is given by $\ell=g^{2}/2M$, with $M$ the BH mass. Nevertheless, as with all the other RBHs considered, here we shall treat the GCSV RBH as a toy model for a regular space-time possessing a Minkowski core.
@@ II.4 Hayward-like black hole
For all of the three RBHs discussed earlier, the temperature decreases with increasing regularizing parameter, while converging to the Schwarzschild value $T_{\text{Sch}}=1/8\pi M$ as $\ell\to 0$, see Fig. 1. It is therefore not unreasonable to expect that evaporation constraints on $f_{\text{pbh}}$ may be weakened if the DM consists of these PRHBs (this will of course be checked explicitly later). We have verified that such a behaviour of temperature versus regularizing parameter holds for most PRBHs of phenomenological interest considered in the literature.  [^3] This is actually not unrelated to the fact that most RBHs, and more generally hairy BH solutions, feature shadows whose size decreases relative to the size of the Schwarzschild BH shadow, i.e. $3\sqrt{3}M$, see Ref. Vagnozzi et al. (2023) for further more detailed discussions (see also Refs. Vertogradov and Övgün (2024a, b); Vertogradov et al. (2024)). In fact, the equations governing BH temperature and shadow size bear some resemblance, so an increase/decrease in one is expected to lead to an increase/decrease in the other. Ref. Vagnozzi et al. (2023) also presents a few space-times whose shadow radii increase with increasing hair parameter, and for these we would expect the temperatures to increase as well. However, for a few of these space-times we explicitly tested (which, we stress, are phenomenological in nature), the Teukolsky equation used to calculate the greybody factors turned out to be numerically very complicated, which is why we opted for proposing the equally phenomenological Hayward-like BH discussed here. For phenomenological reasons, especially in light of our expectation (verified a posteriori) that such a behaviour would lead to weaker bounds on $f_{\text{pbh}}$ and thereby the asteroid window further opening, it could therefore be desirable to also contemplate a case where the BH temperature actually increases with increasing regularizing parameter.
@@ III.1 Greybody factors
For the static, spherically and tr-symmetric metrics given by Eq. (4) which we consider, the Teukolsky equation in spherical coordinates is separable. A key role in computing the GBFs is played by the radial Teukolsky equation, which we now report in full generality for the class of metrics in question. Using the Newman-Penrose (NP) formalism Newman and Penrose (1962), the Teukolsky equation governing the evolution of (massless) perturbations of different spin can be condensed into a single master equation Teukolsky (1973):
@@
$\begin{split}&f_{\text{B}}(x)=1-\frac{(1+\ell^{2})^{3/2}(x+1)^{2}}{\left(\ell^{2}+(x+1)^{2}\right)^{\frac{3}{2}}}\,,\\ &f_{\text{H}}(x)=1-\frac{(x+1)^{2}}{(1-\ell^{2})\left((x+1)^{3}-\frac{\ell^{2}}{\ell^{2}-1}\right)}\,,\\ &f_{\text{GCSV}}(x)=1-\frac{e^{\ell-\frac{\ell}{x+1}}}{x+1}\,,\\ &f_{\text{H-l}}(x)=1-\frac{(1+\ell-\ell^{2})(x+1)^{2}}{\left((x+1)^{3}+\ell-\ell^{2}(x+1)\right)}\,,\end{split}$ (24)
for the Bardeen, Hayward, GCSV, and Hayward-like space-times respectively.
@@
$\delta_{s}=\tau\frac{ie^{i\pi s}(2\omega)^{2s-1}\Gamma\left(1-s-\frac{2i\omega}{\tau}\right)}{\Gamma\left(s-\frac{2i\omega}{\tau}\right)}$ (28)
Using the method discussed above, we compute the GBFs for perturbations of different spin on the backgrounds of the four RBH space-times discussed earlier, for different values of the regularizing parameter $\ell$. In the specific case $s=1$, we have checked that calculating the GBFs up to $l=4$ is sufficient for our purposes. The GBFs we calculate are then used to characterize the Hawking evaporation spectra, as we will discuss shortly.
@@ III.2 Evaporation spectra
Figure 4: As in Fig. 2, but for Ghosh-Culetu-Simpson-Visser black holes, with identical values of the regularizing parameter $\ell/r_{H}$ and identical color coding.
@@
Contrarily to what we observe for these three RBHs, we see that the intensity of the spectra for Hayward-like RBHs substantially increase with increasing $\ell$ for sufficiently large energies $E\gtrsim{\cal O}({\text{MeV}})$, whereas for lower energies they very slightly decrease. This behaviour agrees qualitatively with the expectations which motivated us to construct the Hayward-like RBH in first place, and with the temperature evolution shown in Fig. 1. In addition, we observe that an increase in the regularizing parameter leads to a much richer structure in the spectra: the observed multi-peak structure is the result of the contribution of different $l$ modes in the emission being disentangled more clearly. A similar behaviour is actually observed in the Kerr BHs evaporating spectra which, by virtue of their rotation, lead to a more complex spectrum where the contribution of high-$l$ modes, while still subdominant with respect to the $l=1$ one, emerges more clearly (although we have explicitly checked that adding $l>4$ modes has virtually no effect on our results). In the case of Hayward-like RBHs, we remain agnostic as to the physical reason why such a structure is observed, and leave a more detailed study to follow-up work.
@@ III.3 Evaporation constraints
We work under the commonly adopted assumption that PBHs are isotropically distributed on sufficiently large scales. Therefore, the flux resulting from their evaporation and reaching us today is given by the redshifted sum of the contributions from all evaporating PBHs in our Universe, and can be used to constrain the average extragalactic distribution of DM in the form of PBHs. We also work within the (also commonly adopted) approximation of monochromatic mass distributions (which can be expected if the formation mechanism arises from an amplification of the power spectrum at a very specific scale), although the effect of extended mass distributions is the subject of active research Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024). Finally, as discussed earlier, we only consider the primary photon contribution, as the secondary component resulting from the decay into photons of other unstable particles is verified a posteriori to be negligible given the mass range of interest. While all these are clearly approximations, albeit widely adopted ones, we are confident that they are appropriate given the aim of our work. Our main goal is to examine how the limits on $f_{\text{pbh}}$ change when moving from the Schwarzschild PBH framework to that of the regular metrics presented in Sec. II, potentially opening or closing the asteroid mass window. It is more than reasonable to expect that the shift in constraints relative to the Schwarzschild case $\delta f_{\text{pbh}}$ is only weakly affected by the above approximations. In other words we expect these approximations to have similar impacts on the contraints on $f_{\text{pbh}}$ relative to the Schwarzschild BHs and the RBHs discussed in Sec. II, therefore leading to negligible effects on the shift $\delta f_{\text{pbh}}$, in which we are ultimately interested. In any case, such approximations also allow for a more direct comparison to several previous works and therefore we consider them appropriate for our pilot study, but their impact should definitely be explored in follow-up works. Finally, note that we are tacitly assuming that PBHs cluster in the galactic halo in the same way as other forms of DM (unless they are extremely large, which is not the case for the mass range of interest).
@@
We assume a spatially flat $\Lambda$CDM cosmological model in specifying the expansion rate entering into Eq. (31), with the same cosmological parameters as in Ref. Carr et al. (2010). This allows us to robustly cross-check our Schwarzschild constraints on $f_{\text{pbh}}$ against those reported in the seminal Ref. Carr et al. (2010), although we stress that our constraints are very stable against reasonable changes in the values of the cosmological parameters. Once the cosmological model is fixed, all the relevant quantities in Eq. (31) are known except for the present-day PBH number density, $n_{\text{pbh}}(t_{0})$, which can be constrained from EGRB observations and is ultimately related to $f_{\text{pbh}}$. More specifically, for any given value of the PBH mass $M_{\text{pbh}}$, through Eqs. (29,31,32) we can compute the unnormalized photon flux $I(E_{\gamma 0})/n_{\text{pbh}}(E_{\gamma 0})$, and adjust the normalization $n_{\text{pbh}}(E_{\gamma 0})$ by comparing against EGRB observations (as we will explain shortly). This procedures gives us an upper limit on $n_{\text{pbh}}(E_{\gamma 0})$, which can be translated into an upper limit on $f_{\text{pbh}}$ as follows:

### v2

@@ II Regular black holes
Figure 1: Evolution of the temperatures (normalized to the temperature of Schwarzschild black holes, $T_{\text{Sch}}=1/8\pi M$) as a function of the regularizing parameter $\ell$ (normalized to the event horizon radius $r_{H}$) for the four regular black holes studied in this work: the Bardeen (blue solid curve, Sec. II.1), Hayward (red dashed curve, Sec. II.2), Culetu-Ghosh-Simpson-Visser (green dotted curve, Sec. II.3), and Hayward-like (magenta dash-dotted curve) regular space-times. In the first three cases the temperature is a monotonically decreasing function of $\ell$, whereas in the last case it increases monotonically with $\ell$ after an initial dip, with important phenomenological consequences resulting from these different behaviours. Note that the range of allowed values of $\ell/r_{H}$ is different for the four regular black holes.
@@
A final caveat is in order before discussing the RBH metrics we consider. The latter are all regular in the sense of having finite curvature invariants $R\equiv g^{\mu\nu}R_{\mu\nu}$, $R_{\mu\nu}R^{\mu\nu}$, and ${\cal K}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. However, a more stringent criterion for regularity is that of geodesic completeness, which does not necessarily imply finiteness of curvature invariants and viceversa. A number of “popular” RBHs have indeed been shown to have finite curvature invariants but to be geodesically incomplete Zhou and Modesto (2023). This includes the well-known Bardeen and Hayward RBHs, which are among the ones we shall consider here. Nevertheless, given the significant interest in these metrics, the fact that they are widely taken as prototypes for RBHs, and our phenomenological goal of going beyond Schwarzschild PBHs, we will take these space-times into consideration, while cautioning the reader about the above issues, and therefore that these metrics should be considered nothing more than phenomenological toy models at this stage. Note, in addition, that the stability of RBHs featuring inner horizons is currently a matter of debate in the literature Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023).
@@ II.1 Bardeen black hole
The Bardeen BH is easily one of the best known RBHs, and one of the first ones to ever have been proposed Bardeen (1968). It is characterized by the following metric function:  [^2] Note that in all the metrics considered, the parameter $M$ appearing in the metric function can always be unambiguously identified with the BH mass (either the Komar, ADM, Misner-Sharp-Hernandez, or Brown-York mass). This is important as the constraints we subsequently derive on $f_{\text{pbh}}$ as a function of $M_{\text{pbh}}$ will implicitly identify the latter with $M$.
@@ II.2 Hayward black hole
Another widely known regular space-time is the Hayward RBH Hayward (2006), characterized by the following metric function:
@@
If expressed in terms of BH mass $M$, the regularizing parameter for the Hayward BH is subject to the same limit as that of the Bardeen BH, i.e. $\ell\leq\sqrt{16/27}\,,M$. On the other hand, if expressed in terms of the more physically motivated horizon radius, the limit is instead $\ell\lesssim 0.57\,r_{H}$. We note that the Schwarzschild metric function is recovered in the $\ell\to 0$ limit.
@@
Just as the Bardeen RBH possesses a dS core, so does the Hayward RBH. Indeed, introducing a dS core characterized by a (positive) cosmological constant $\Lambda=3/\ell^{2}$ in order to prevent the central singularity was precisely the original justification for the Hayward BH which, just like its Bardeen counterpart, was proposed on purely phenomenological grounds. Nevertheless, potential theoretical origins for the Hayward BH have been investigated, and range from corrections to the equation of state of matter at high density Sakharov (1966); Gliner (1966), finite density and finite curvature proposals Markov (1982, 1987); Mukhanov and Brandenberger (1992), theories of non-linear electrodynamics Kumar et al. (2020b); Kruglov (2021), and more generally as the result of corrections due to quantum gravity Addazi et al. (2022); Alves Batista et al. (2023). Just as with the Bardeen RBH, we shall here treat the Hayward RBH as a model-agnostic phenomenological toy model for a singularity-free space-time.
II.3 Culetu-Ghosh-Simpson-Visser black hole
@@
The regular space-times considered so far featured dS cores, which in itself is a very common feature of several RBH metrics. Nevertheless, another interesting phenomenological possibility consists in considering “hollow” RBHs wherein the central singularity is replaced by an asymptotically Minkowski core, where the associated energy density and pressure asymptote to zero. This is quite unlike the case of the dS core where the energy density asymptotes to a finite value associated to a positive cosmological constant, and the pressure asymptotes to an equal but opposite value. Possible theoretical/mathematical motivations for considering RBHs with Minkowski cores include the fact that the vanishing energy density can significantly simplify the physics in the deep core, whereas the otherwise messy solutions to polynomial equations (which often cannot be written down in closed form) can be traded for arguably more elegant special functions, resulting in the space-time being more tractable. Our physical motivation in considering this class of BHs is instead to broaden the range of physical properties and phenomenological implications of PRBHs, going beyond the dS core RBHs studied thus far.
@@
With this in mind, we consider a RBH featuring a Minkowski core, independently studied by Culetu Culetu (2013, 2015), Ghosh Ghosh (2015), as well as Simpson and Visser Simpson and Visser (2019). Although such a RBH does not have any particular name associated to it in the literature, here we refer to it as CGSV BH (from the initials of the four authors above). The space-time is characterized by the following metric function:
$f_{\text{CGSV}}(r)=1-\frac{2M}{r}\exp\left(-\frac{\ell}{r}\right)\,.$ (10)
@@
The horizon radius $r_{H}$, for which a closed form expression is not available in the Bardeen and Hayward cases, here is given by:
@@
or, alternatively, $0\leq\ell<r_{H}$. While the CGSV BH was original introduced purely on phenomenological/mathematical grounds, it was shown in Refs. Kumar and Ghosh (2021); Singh et al. (2022) that such a space-time can emerge within the context of GR coupled to a specific non-linear electrodynamics source. In this case, denoting by $g$ the non-linear electrodynamics coupling constant/charge, the regularizing parameter $\ell$ is given by $\ell=g^{2}/2M$, with $M$ the BH mass. Nevertheless, as with all the other RBHs considered, here we shall treat the CGSV RBH as a toy model for a regular space-time possessing a Minkowski core.
@@ II.4 Hayward-like black hole
For all of the three RBHs discussed earlier, the temperature decreases with increasing regularizing parameter, while converging to the Schwarzschild value $T_{\text{Sch}}=1/8\pi M$ as $\ell\to 0$, see Fig. 1. It is therefore not unreasonable to expect that evaporation constraints on $f_{\text{pbh}}$ may be weakened if DM consists of these PRBHs (this will of course be checked explicitly later). We have verified that such a behaviour of temperature versus regularizing parameter holds for most PRBHs of phenomenological interest considered in the literature.  [^3] This is actually not unrelated to the fact that most RBHs, and more generally hairy BH solutions, feature shadows whose size decreases relative to the size of the Schwarzschild BH shadow, i.e. $3\sqrt{3}M$, see Ref. Vagnozzi et al. (2023) for further more detailed discussions (see also Refs. Vertogradov and Övgün (2024a, b); Vertogradov et al. (2024)). In fact, the equations governing BH temperature and shadow size bear some resemblance, so an increase/decrease in one is expected to lead to an increase/decrease in the other. Ref. Vagnozzi et al. (2023) also presents a few space-times whose shadow radii increase with increasing hair parameter, and for these we would expect the temperatures to increase as well. However, for a few of these space-times we explicitly tested (which, we stress, are phenomenological in nature), the Teukolsky equation used to calculate the greybody factors turned out to be numerically very complicated, which is why we opted for proposing the equally phenomenological Hayward-like BH discussed here. For phenomenological reasons, especially in light of our expectation (verified a posteriori) that such a behaviour would lead to weaker bounds on $f_{\text{pbh}}$ and thereby the asteroid mass window further opening, it could therefore be desirable to also contemplate a case where the BH temperature actually increases with increasing regularizing parameter.
@@ III.1 Greybody factors
For the static, spherically and tr-symmetric metrics given by Eq. (4) which we consider, the Teukolsky equation in spherical coordinates is separable. A key role in computing the GBFs is played by the radial Teukolsky equation, which we now report in full generality for the class of metrics in question. Using the Newman-Penrose (NP) formalism, the Teukolsky equation governing the evolution of (massless) perturbations of different spin can be condensed into a single master equation Teukolsky (1973):
@@
$\begin{split}&f_{\text{B}}(x)=1-\frac{(1+\ell^{2})^{3/2}(x+1)^{2}}{\left(\ell^{2}+(x+1)^{2}\right)^{\frac{3}{2}}}\,,\\ &f_{\text{H}}(x)=1-\frac{(x+1)^{2}}{(1-\ell^{2})\left((x+1)^{3}-\frac{\ell^{2}}{\ell^{2}-1}\right)}\,,\\ &f_{\text{CGSV}}(x)=1-\frac{e^{\ell-\frac{\ell}{x+1}}}{x+1}\,,\\ &f_{\text{H-l}}(x)=1-\frac{(1+\ell-\ell^{2})(x+1)^{2}}{\left((x+1)^{3}+\ell-\ell^{2}(x+1)\right)}\,,\end{split}$ (24)
for the Bardeen, Hayward, CGSV, and Hayward-like space-times respectively.
@@
$\delta_{s}=\tau\frac{ie^{i\pi s}(2\omega)^{2s-1}\Gamma\left(1-s-\frac{2i\omega}{\tau}\right)}{\Gamma\left(s-\frac{2i\omega}{\tau}\right)}\,,$ (28)
where $\Gamma$ is the $\Gamma$ function. Using the method discussed above, we compute the GBFs for perturbations of different spin on the backgrounds of the four regular metrics discussed earlier, for different values of the regularizing parameter $\ell$. In the specific case $s=1$, we have checked that calculating the GBFs up to $l=4$ is sufficient for our purposes. The GBFs we calculate are then used to characterize the Hawking evaporation spectra.
@@ III.2 Evaporation spectra
Figure 4: As in Fig. 2, but for Culetu-Ghosh-Simpson-Visser black holes, with identical values of the regularizing parameter $\ell/r_{H}$ and identical color coding.
@@
Contrarily to what we observe for these three PRBHs, we see that the intensity of the spectra for Hayward-like PRBHs substantially increase with increasing $\ell$ for sufficiently large energies $E\gtrsim{\cal O}({\text{MeV}})$, whereas for lower energies they very slightly decrease. This behaviour agrees qualitatively with the expectations which motivated us to construct the Hayward-like RBH in first place, and with the temperature evolution shown in Fig. 1. In addition, we observe that an increase in the regularizing parameter leads to a much richer structure in the spectra: the observed multi-peak structure is the result of the contribution of different $l$ modes in the emission being disentangled more clearly. A similar behaviour is actually observed in the evaporation spectra of Kerr BHs which, by virtue of their rotation, lead to a more complex spectrum where the contribution of high-$l$ modes, while subdominant with respect to the $l=1$ mode, emerges more clearly (although we have explicitly checked that adding $l>4$ modes has virtually no effect on our results). In the case of Hayward-like RBHs, we remain agnostic as to the physical reason why such a structure is observed, and leave a more detailed study to follow-up work.
@@ III.3 Evaporation constraints
We work under the commonly adopted assumption that PBHs are isotropically distributed on sufficiently large scales. Therefore, the flux resulting from their evaporation and reaching us today is given by the redshifted sum of the contributions from all evaporating PBHs in our Universe, and can be used to constrain the average extragalactic distribution of DM in the form of PBHs. Furthermore, we work within the (also commonly adopted) approximation of monochromatic mass distributions (which can be expected if the formation mechanism arises from an amplification of the power spectrum at a very specific scale), although the effect of extended mass distributions is the subject of active research Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b). Finally, as discussed earlier, we only consider the primary photon contribution, as the secondary component resulting from the decay into photons of other unstable particles is verified a posteriori to be negligible given the mass range of interest. While all these are clearly approximations, albeit widely adopted ones, we are confident that they are appropriate given the aim of our work. Our main goal is in fact to examine how the limits on $f_{\text{pbh}}$ change when moving from the Schwarzschild PBH framework to that of the regular metrics presented in Sec. II, potentially opening or closing the asteroid mass window. It is more than reasonable to expect that the shift in constraints relative to the Schwarzschild case, $\delta f_{\text{pbh}}$, is only weakly affected by the above approximations. In other words we expect these approximations to have similar impacts on the constraints on $f_{\text{pbh}}$ relative to Schwarzschild BHs for the different RBHs discussed in Sec. II, therefore leading to negligible effects on the relative shift $\delta f_{\text{pbh}}$, which is the quantity we are ultimately interested in. At any rate, the adopted approximations also allow for a more direct comparison to several previous works, and we therefore consider them appropriate for our pilot study, while stressing that their impact should definitely be explored in future follow-up works. Finally, note that we are tacitly assuming that PBHs cluster in the galactic halo in the same way as other forms of DM (unless they are extremely large, which is not the case for the mass range of interest).
@@
We assume a spatially flat $\Lambda$CDM cosmological model in specifying the expansion rate entering into Eq. (31), with the same cosmological parameters as in Ref. Carr et al. (2010). This allows us to robustly cross-check our Schwarzschild constraints on $f_{\text{pbh}}$ against those reported in the seminal Ref. Carr et al. (2010), although we stress that our constraints are very stable against reasonable changes in the values of the cosmological parameters. Once the cosmological model is fixed, all the relevant quantities in Eq. (31) are known except for the present-day PBH number density, $n_{\text{pbh}}(t_{0})$, which can be constrained from EGRB observations and is ultimately related to $f_{\text{pbh}}$. More specifically, for any given value of the PBH mass $M_{\text{pbh}}$, through Eqs. (29,31,32) we can compute the unnormalized photon flux $I(E_{\gamma 0})/n_{\text{pbh}}(t_{0})$, and adjust the normalization $n_{\text{pbh}}(t_{0})$ by comparing against EGRB observations (as we will explain shortly). This procedures gives us an upper limit on $n_{\text{pbh}}(t_{0})$, which can be translated into an upper limit on $f_{\text{pbh}}$ as follows: